9=13x-7-5x what does x equal show work thanks!
Found 2 solutions by funmath, AnlytcPhil:
Answer by funmath(2933)

9=13x-7-5x
9=(13-5)x-7
9=8x-7
9+7=8x-7+7
16=8x
16/8=8x/8
2=x
Check your answer by substituting your answer into the equation to see if it works.
9=13(2)-7-5(2)
9=26-7-10
9=26-17
9=9 Yeah! We're right!
x=2

Answer by AnlytcPhil(1806)


9=13x-7-5x what does x equal show work thanks!


          9 = 13x - 7 - 5x

First combine the like terms on the right which are
13x and -5x.  That gives you 8x. So you take out those 
two terms and put in the one term "8x". So now you 
have this:

          9 = 8x - 7 

Since you want to get the x-term, 8x, all by itself
we get rid of the -7 term on the right by adding its
opposite, +7, to both sides

          9 = 8x - 7
         +7       +7
        -------------
         16 = 8x

When we get to this stage, where there is only one 
term on one side consisting of a coefficient and the
unknown, then you divide both sides by that 
coefficient, which is 8.
 
         16     8x
        ---- = ----
          8      8

the 8 goes into the 16 2 times on left, and the 8's
cancel on the right:

          2     1
         16     8x
        ---- = ----
          8      8
          1      1

so the answer is:

           2 = x

Edwin